Agenda/Minutes on 2021/2/12
CRY+VIS Interface Meeting
Participants: Takahashi, Sato, Aso, Hirata, Ikeda, Ushiba, Miyo, Tomaru, Hagiwara
- Schedule for IX: Release is scheduled on 2/18.
- Schedule for IY: Disconnection is scheduled on 3/11.
- Schedule for EX: Connection is scheduled in the week of 3/15.
- Corrugate tube fixing jig: Re-measurement is planed in 2/26-3/1 at IYV.
- Commissioning in air: Commissioning of IX will be started. Other schedule has not been decided yet.
- Commissioning in vacuum: Closing of DN800 flange is available within a half of day.
- Acceptance check: Cryo-payload in IX and Type-A tower in EY are on going.
- PF actuator: Not evaluated in IX, but after the commissioning for suspension.
- Jig for FC painting: Delivered.
Participants: Takahashi, Fabian, Hirata, Ikeda, Miyo, Sato, Aso, Sugiyama, Ishizaki, Tanaka, Washimi
Ikeda-san worked on following items;
- He measured the length of the network cable (15956), and now is preparing to order all cables we need with DGS group.
- He got message from manufacturer that it is difficult to operate the picomotor for long time more 24 hours. (15958). So he is now considering to request AEL group to make the circuit box that turns power on and off with BIO signal like the stepper motor circuit.
- He made the new MEDM screen for stepper motor on MN stage which is developed by Nishimoto-kun. We can open the screen from [sitemap]/[STEPPINGMOTOR]/[CRY_OVERVIEW] (15965).
15737 We investigated the network connection. , satoru.ikeda
15956 Comment, satoru.ikeda
15958 About Picomotor's 24-hour operation , satoru.ikeda
15965 Integration of Cryo Payload Phytron motor driver into MEDM , satoru.ikeda
15983 I posted this by mistake. , satoru.ikeda
Setup of the F2 stage in EY: installation of suspension wire, cabling, calibration of LVDT, check of FR range, tuning
15950 Setup of F2 , ryutaro.takahashi
15982 Setup of F2 , ryutaro.takahashi
Thursday 15993 Setup of F2 , ryutaro.takahashi
15959 Connection between F3 and F2 , naohisa.satou
15972 Cabling between F3 and F2 , naohisa.satou
Connection of the cryo-payload in IX, mounted the geophones onto the IP in IX, and compressed the F0 blades in IX.
15968 Connection to cryo-payload , ryutaro.takahashi
15974 Comment, ryutaro.takahashi
15968 Connection to cryo-payload , ryutaro.takahashi
15981 Setup of pre-isolator , ryutaro.takahashi
Akutsu-san found the optical lever light hits on the edge of the MCE mirror during his investigation about IMC ghost beams. We need to re-schedule to work on IMC suspension for this problem and also other issues reported in 15916. Please discuss it later.
15987 MCe oplev , tomotada.akutsu
(1) Miyo fixed the swapped cabling on the BF-damper for PRM with Nakagaki-san, and after that, the noise coupling of the transfer function at higher frequency is decreased to the same behavior of other PR suspensions (15994). By the way, the same noise coupling was reported in Type-A suspension 12678. Is it caused by the same error?
(2) Ikeda-san touched the picomotor of PRM for his first operation (15957) before measure the calibration factor between picomotor and OSEM.
9449 PRM glitch investigation: , koki.okutomi
15957 Checking the operation of PRM's Picomotor PRM_BF and PRM_IM , satoru.ikeda
Ikeda-san tested the picomotor operation and he did rough calibration for PR3-IM picomotor (15977). During the test for BF picomotor, it was stucked (15976). So we need open the top chamber to fix that.
15977 PR3 IM PITCH, ROLL, YAW's Picomotor is operated to measure the travel and drive range , satoru.ikeda
15976 Picomotor of PR3_BF_T(Motor2) got stuck , satoru.ikeda
15979 Comment, satoru.ikeda
SR3 ballast mass adjustment highlights
On top of Intermediate Mass
- We recovered the oplev in order to use is as a reference for pitch.
Aim for the BF LVDT: 311 um per buoyancy correction.
- After locking the IRM we checked the IM was free to move.
- After we locked the payload and BF to the security structure, we removed the H1 OSEM flag and then the front and back panels of the IRM.
In two steps we removed a total of 100 grams from the IM and got the BF LVDT to 308 um at the temperature of the moment.
- Once we finished the adjustment we assembled the IRM and put the OSEM flag back with magnet with the correct polarity.
Despite we used the oplev as a reference, apparently we disturbed the pitch by 496 urad approximately. This is should be within picomotor range. Yet to be adjusted.
We measured five IM transfer functions and the system looks very healthy. We didn't measure IM-V because the top filter is locked.
We measured the position of the IM-P picomotor.
On top of Bottom Filter (klog yet to be written)
F1 LVDT aim 416 um per buoyancy effect.
- From IM ballast mass adjustment we calculated how much mass we require for the BF: 2 * 45 grams.
We quickly tested the calculation this morning with temporary ballast masses and got it to 412 um.
- On Monday we will put the masses properly and fix them.
15928 Oplev installation and alignment recovery. , fabian.arellano
15953 Comment, fabian.arellano
15963 IM ballast mass adjustment , fabian.arellano
- Measured the TFs of old and new combiner ciruits. 30dB attenuation was found at 100Hz.
- New combiner circuit was sent to Kamioka.
- Type-B/Bp remedy work is on going.
- Installation of Type-A is on going in EY.
- Procedure documents.
- Criteria in the acceptance test.
[Ikeda] How to rescue the stucked picomotor named BF-T for PR3. klog#15976
[Miyo] When and who work on the IMC repairing? This work is not included in the VIS schedule.
- Making of extension cables with the connector.
- Fixing of stray wires on the pillar.
- Fixing of the 22-pin Burndy cable to the connector.
[Miyo] Please write klog entry as a comment if it is related work to other entry as Fabian did on klog#15963. If you do so, it is very helpful for future workers to find whether the work is done or not.
The week from 2/15
- Takahashi: 2/15-19
- Sato: 2/15-19
- Hirata: 2/15-19
- Yano: 2/15-19
The week from 2/22
- Takahashi: 2/23-2/28
- Sato: 2/23-26
- Hirata: 2/23-26
- Yano: 2/23-26
- On 2020/2/19(Fri)